# Formatting Instructions For NeurIPS 2025

### v1

@@ 4.1 Quantitative and qualitative comparison
ResNet50 87% / 63% 85% / 60% 87% / 65%
ConvNeXt 92% / 72% 94% / 76% 95% / 80%
SwinV2 94% / 77% 95% / 79% 95% / 81%
Robustness evaluation To assess the robustness of the presented results, we carry out cross-model comparisons where we measure model performance based on samples generated by conditioning on the feature map extract from a different model. The results for this experiment are compiled in Tab. 2. It turns out that all three sets of samples (conditioned on features generated by the three different backbones) transfer quite remarkably to other models. In the supplementary material, we also present results supporting the robustness of our approach when applied to out-of-distribution (OOD) samples.
@@ 4.2 Application: FeatInv-Viz – Visualizing concept steering in input space
Concept steering in input space In generative NLP, steering is sometimes used to verify concept interpretations by reducing or magnifying concepts in the model activations and observing how this changes the generated output text, as famously demonstrated at the example of the Golden Gate concept [5] in Claude 3 Sonnet. This approach is not directly applicable to vision classifiers. However, with our method of inverting model representations from feature to input space, we can observe the effect of concept steering within hidden model activations in the input representation space instead of the output. This enables a novel method for concept visualization, with benefits over existing approaches (see below).
Concept definition Concepts are typically defined as structures in feature space such as individual neurons, single directions or multi-dimensional subspaces. Many concept-based XAI methods define a way to decompose a feature vector into concepts from a dictionary/concept bank [9]. In this work, we use concepts from multi-dimensional concept discovery (MCD) [25], which defines concepts as linear subspaces in feature space. Nevertheless, our approach is applicable to any concept discovery method.
Concept visualization through attenuated feature maps A common challenge for unsupervised concept discovery methods is inferring the meaning of discovered concepts. To address this, we steer a concept in feature space and observe the effect in input space. Specifically, we attenuate coefficients for the concept under consideration to 25% (see the Supplementary Material for a details). Then, we use FeatInv to map the original and the modified feature map to input space using identical random seeds for the diffusion process. By comparing the resulting images, we gain insights into how the concept is expressed in input space. We call this method FeatInv-Viz and present it in Algorithm 1.
@@
Exemplary results Fig. 4 shows exemplary concept steering visualizations for four samples from the Indigo Bunting class. Here, we decomposed ConvNeXt’s feature space into three linear concept subspaces. FeatInv-Viz provides a visualization of these concepts in input space. The method provides a very finegrained visualization of which specific regions in input space change upon steering each concept in feature space.
Benefits We emphasize that FeatInv-Viz extends commonly used concept activation maps in two ways: First, it provides a finegrained visualization rather than a coarse upscaling [2, 25] of a lower-resolution feature map. Second, it goes beyond merely verifying alignment with a predefined concepts [2], by providing counterfactual information from targeted feature-map manipulations.
@@ 4.3 Application: Investigating the composite nature of the feature space
[…]
@@ A.1 Additional models
For the ablation studies, we consider ResNet50B222timm model weights: resnet50.a1_in1k [26] as second ResNet-model to study the impact of the training procedure, in addition to the models studied in the main text.
Choice of conditional encoder Next to the bilinear upsampling followed by a shallow CNN, we explore a different design choice for the conditional encoder: To this end, we explore the use of cross-attention as used in the Perceiver architecture [13], which uses a learnable representation of predefined size and connects it via cross-attention to the representation that serves as input for the mapping. Unlike the previous approach, this approach is not subject to any locality assumptions and therefore the most flexible approach that is in particular suitable for model architectures without built-in locality assumptions such as vision transformers. We carry out a comparison between the two conditional encoder models for the case of the SwinV2 ViT, where we expect the impact to be most pronounced as the ViT operates on visual tokens, which might not align well with the upscaling operation in the convolutional encoder. On the contrary, our results indicates that the convolutional encoder yields better classification results and therefore for the following experiments we restrict ourselves to the convolutional encoder.
Class-conditional baseline We tested the models on an unconditional baseline to see if MiniSD is able to generate good representations of the classes and if they can be correctly classified. To define the classes for the input prompt as precisely as possible, we use the WordNet hierarchy and create the prompt as follows: ‘a high-quality, detailed, and professional image of a CLASS, which is a kind of SUPERCLASS’ as for example ‘a high-quality, detailed, and professional image of a tench, which is a kind of cyprinid’. The unconditional baseline created consists of 10 samples per class, i.e. 10,000 samples in total.
Table 3: Additional model insights: For the different backbones, we indicate the corresponding canonical input sizes as well as the top5(1) accuracy on an unconditional dataset, created by MiniSD. Since only the base model is used for the prediction of these unconditional samples, the results do not differ between unpooled and pooled as well as for the different SwinV2 architectures. We show three performance metrics: Cosine similarity in feature space (cosine-sim), calculated by taking the median of the cosine similarity of all superpixels, top5(1) matches using the top1 prediction of the original sample as ground truth (top5(1) match) and FID-scores (FID) to assess the quality of the generated samples. We consider generative models conditioned on unpooled feature maps (rows 1-5) and models conditioned on pooled feature maps (rows 6-9).
@@
unpooled ResNet50 $224\times 224$ 89% (68%) 0.40 87% (63%) 12.32
ResNet50B 88% (65%) 0.50 90% (68%) 18.58
ConvNeXt $288\times 288$ 92% (71%) 0.61 94% (76%) 7.14
SwinV2 $384\times 384$ 93% (72%) 0.57 95% (81%) 11.56
SwinV2(Perceiver) 0.37 79% (55%) 22.72
pooled ResNet50 $224\times 224$ 89% (68%) 0.09 47% (22%) 26.37
@@
ConvNeXt $288\times 288$ 92% (71%) 0.18 42% (18%) 31.01
SwinV2 $384\times 284$ 93% (72%) 0.14 47% (23%) 36.21
A.2 FID scores vs. Guidance Scale
@@
In Fig. 6, we present several original images alongside their respective reconstructions, generated using the model trained with the ConvNeXt backbone. These samples were produced with varying levels of guidance scale. Since the guidance scale significantly influences the output, we evaluated the model’s performance across different scale settings, again using the ConvNeXt-based model. To quantify the quality of the generated images, we employed FID and cosine distance, as illustrated in Fig. 7. We opted to use cosine distance (rather than similarity) to ensure that, in both metrics, lower values indicate better performance, simplifying comparison. This suggests that the unpooled feature map requires minimal text guidance, whereas for the pooled variant, guidance is of substantial importance, as natural looking images cannot be generated without its influence.
@@
Figure 6: Unpooled and pooled reconstructions with different guidance scales from ConvNeXt feature maps. The cosine similarity between the original feature map and that of the reconstruction is noted at the bottom edge of the images. While increasing the guidance scale in unpooled samples results in higher color saturation and more distorted object shapes, the same increase improves object fidelity in pooled feature maps. This trend is also roughly reflected in the corresponding cosine distances.
00111122223333444455556666777788889999101010101111111112121212131313130020202020404040406060606080808080guidance scaleFID scoreunpooled FIDpooled FID0.40.40.40.40.60.60.60.60.80.80.80.81111cosine distanceunpooled cos-distpooled cos-dist
Figure 7: Impact of guidance scale on generation quality (FID) and semantic alignment (cos-dist). FID values (solid lines) and cosine distance (dashed lines) for models with ConvNeXt backbone with and without pooling. Lower values in both metrics indicate better performance. The unpooled model (blue) achieves optimal quality at a guidance scale of 1.75, with both low FID and low cosine distance. In contrast, the pooled model (red) performs significantly worse at all scales, with increasing FID and cosine distance as the guidance scale decreases. This model reaches its best performance near a scale of 8.0.
A.3 Generation on unpooled vs. pooled feature maps
@@
Appendix B OOD evaluation
In Fig. 9, we qualitatively test the robustness of our findings by conditioning on samples that are slightly out of the model scope of the original models finetuned on ImageNet. To this end, we use the BAID [27] dataset, which differs in style from the samples in ImageNet. The ControlNet trained on ConvNeXt features still shows a good reconstruction quality of the semantic content but the style of the reconstruction and the original image differ more strongly than in the in-distribution case in Fig. 3. Nevertheless, the results speak for the robustness of the proposed approach.
Figure 9: Reconstruction on OOD samples. Comparison of randomly selected original and generated samples from the BAID [27] dataset, which differs in style from the samples in ImageNet. The basis for the reconstruction was the ControlNet trained on ConvNeXt features, which received the ConvNeXt features of the samples shown as input. The cosine similarity between the original feature map and that of the reconstruction is noted at the bottom edge of the images
Appendix C Details on concept steering

### v2

@@ 4.2 Robustness evaluation
ResNet50 91% / 70% 90% / 68% 91% / 71%
ConvNeXt 92% / 73% 94% / 77% 96% / 81%
SwinV2 94% / 77% 94% / 78% 95% / 80%
Evaluation with OOD samples In Fig. 5, we qualitatively test the robustness of our findings by conditioning on samples that are slightly out of the model scope of the original models finetuned on ImageNet. To this end, we use the BAID (Yi et al., 2023) dataset, which differs in style from the samples in ImageNet. We reconstructed 10,000 images from this dataset. The ControlNet trained on ConvNeXt features still shows a good reconstruction quality of the semantic content but the style of the reconstruction and the original image differ more strongly than in the in-distribution case in Fig. 3. The average cosine similarity is 0.48 and the FID value is 12.98. Nevertheless, the results speak for the robustness of the proposed approach.
Figure 5: Reconstruction of OOD samples. Comparison of randomly selected original and generated samples from the BAID (Yi et al., 2023) dataset, which differs in style from the samples in ImageNet. The basis for the reconstruction was the ControlNet trained on ConvNeXt features, which received the ConvNeXt features of the samples shown as input. The cosine similarity between the original feature map and that of the reconstruction is noted at the bottom edge of the images
Evaluation with adversarial images We tested FeatInv with adversarial images, which are often misclassified by models trained on ImageNet. For this, we used the ImageNet-A (Hendrycks et al. (2019)) dataset, which contains 7,500 samples. In the case of misclassification, the model can be misled by image characteristics such as small objects or confusing patterns. We reconstructed all images in this dataset using our FeatInv model trained on ImageNet ConNeXt feature maps. The average cosine similarity is 0.56 and the FID score is 12.44. FeatInv is therefore able to reconstruct these unknown difficult images to a good degree. The top5(1) classification match is 0.71(0.38), which means that more than a third of the images are classified as the original image, suggesting that parts of the image relevant for classification are reconstructed in a meaningful way. Nevertheless, these matching values are smaller than those of the reconstructed simple images, indicating that the adversarial images are more complex and difficult to reconstruct.
Evaluation with corrupted images We also tested FeatInv with corrupted images. To this end, we used the ImageNet-C dataset (Hendrycks & Dietterich (2019)), which contains various categories of image corruption of varying degrees of intensity. Here, we focused on the weakest (level 1) and the strongest (level 5) corruption levels. In the supplementary material, we present the FID scores for the respective categories and the respective cosine similarity values as well as both the top 5(1) accuracy matches, which correspond to the cosine similarity results, and qualitative results (see Supplementary Material C). As expected, all corruptions have a negative effect on both reconstruction quality and sample quality. Noise and contrast in particular have a high negative impact on the scores and cannot be reproduced well by FeatInv. However, it is worth noting that level 5 corruption represent represent a significant distortion of the original images and therefore represents a challenging robustness scenario.
@@ 4.3 Application: FeatInv-Viz – Visualizing concept steering in input space
Concept steering in input space In generative NLP, steering is sometimes used to verify concept interpretations by reducing or magnifying concepts in the model activations and observing how this changes the generated output text, as famously demonstrated at the example of the Golden Gate concept (Bricken et al., 2024) in Claude 3 Sonnet. This approach is not directly applicable to vision classifiers. However, with our method of inverting model representations from feature to input space, we can observe the effect of concept steering within hidden model activations in the input representation space instead of the output. This enables a novel method for concept visualization, with benefits over existing approaches (see below).
Concept definition Concepts are typically defined as structures in feature space such as individual neurons, single directions or multi-dimensional subspaces. Many concept-based XAI methods define a way to decompose a feature vector into concepts from a dictionary/concept bank (Fel et al., 2023). In this work, we use concepts from multi-dimensional concept discovery (MCD) (Vielhaben et al., 2023), which defines concepts as linear subspaces in feature space. Nevertheless, our approach is applicable to any concept discovery method.
Concept visualization through attenuated feature maps A common challenge for unsupervised concept discovery methods is inferring the meaning of discovered concepts. To address this, we steer a concept in feature space and observe the effect in input space. Specifically, we attenuate coefficients for the concept under consideration to 25%, see the Supplementary Material D for details. Then, we use FeatInv to map the original and the modified feature map to input space using identical random seeds for the diffusion process. By comparing the resulting images, we gain insights into how the concept is expressed in input space. We call this method FeatInv-Viz and present it in Algorithm 1.
@@
Exemplary results Fig. 6 shows exemplary concept steering visualizations for four samples from the Indigo Bunting class. Here, we decomposed ConvNeXt’s feature space into three linear concept subspaces. FeatInv-Viz provides a visualization of these concepts in input space. The method provides a very finegrained visualization of which specific regions in input space change upon steering each concept in feature space. More examples can be found in the Supplementary Material D.1.
Benefits We emphasize that FeatInv-Viz extends commonly used concept activation maps in two ways: First, it provides a finegrained visualization rather than a coarse upscaling Bau et al. (2017); Vielhaben et al. (2023) of a lower-resolution feature map. Second, it goes beyond merely verifying alignment with a predefined concepts Bau et al. (2017), by providing counterfactual information from targeted feature-map manipulations.
@@ 4.4 Application: Investigating the composite nature of the feature space
In NLP, well-known examples of feature-space arithmetic – e.g. $\text{king}-\text{man}+\text{woman}=\text{queen}$ Mikolov et al. (2013) – have shaped our understanding of embedding geometries. FeatInv offers insights into the composite nature of the feature space in vision models by conditioning on feature maps from two samples. In particular, we investigate the effect of convex linear superpositions of two feature maps. To this end we linearly interpolate between the feature representations of two input samples and visualize reconstructions for different weighted combinations, as shown in Fig. 7. We also indicate the cosine similarity between the reconstruction and the weighted feature map, which is highest for the original feature maps and typically reaches its lowest value for the equally weighted interpolated feature map. This can be seen as an indication that the weighted average of two feature maps is in general not a well-defined operation. Nevertheless, foreground objects from one image and background from a second, seem to be reasonably combined through linear superposition (see e.g. bird, landscape).
@@ 5 Summary and Discussion
Acknowledgements
This work was supported by the Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR) under Project INGVER, grant number 01KX2419.
@@ A.1 Additional models
For the ablation studies, we consider ResNet50B222timm model weights: resnet50.a1_in1k (Wightman et al., 2021) as second ResNet-model to study the impact of the training procedure, in addition to the models studied in the main text.
Choice of conditional encoder Next to the bilinear upsampling followed by a shallow CNN, we explore a different design choice for the conditional encoder: To this end, we explore the use of cross-attention as used in the Perceiver architecture (Jaegle et al., 2021), which uses a learnable representation of predefined size and connects it via cross-attention to the representation that serves as input for the mapping. Unlike the previous approach, this approach is not subject to any locality assumptions and therefore the most flexible approach that is in particular suitable for model architectures without built-in locality assumptions such as vision transformers. We carry out a comparison between the two conditional encoder models for the case of the SwinV2 ViT, where we expect the impact to be most pronounced as the ViT operates on visual tokens, which might not align well with the upscaling operation in the convolutional encoder. On the contrary, our results indicates that the convolutional encoder yields better classification results and therefore for the following experiments we restrict ourselves to the convolutional encoder.
Class-conditional baseline We tested the models on an unconditional baseline to see if MiniSD is able to generate good representations of the classes and if they can be correctly classified. To define the classes for the input prompt as precisely as possible, we use the WordNet hierarchy and create the prompt as follows: ‘a high-quality, detailed, and professional image of a CLASS, which is a kind of SUPERCLASS’ as for example ‘a high-quality, detailed, and professional image of a tench, which is a kind of cyprinid’. The unconditional baseline created consists of 10 samples per class, i.e. 10.000 samples in total.
Table 3: Additional model insights: For the different backbones, we indicate the corresponding canonical input sizes as well as the top5(1) accuracy on an unconditional dataset, created by MiniSD. Since only the base model is used for the prediction of these unconditional samples, the results do not differ between unpooled and pooled as well as for the different SwinV2 architectures. We show three performance metrics: Cosine similarity in feature space (cosine-sim), calculated by taking the mean of the cosine similarity of all superpixels, top5(1) matches using the top1 prediction of the original sample as ground truth (top5(1) match) and FID-scores (FID) to assess the quality of the generated samples. We consider generative models conditioned on unpooled feature maps (rows 1-5) and models conditioned on pooled feature maps (rows 6-9). While reconstructions were generated using model-specific values for control strength and guidance scale, we also observed that good results can be achieved with uniform values. A control strength of 1.5(1) and a guidance scale of 8(1.75) for (un)pooled conditional input seemed to work well for this scenario.
@@
unpooled ResNet50 $224\times 224$ 89% (68%) 0.46 91% (70%) 11.49
ResNet50B 88% (65%) 0.50 90% (70%) 11.51
ConvNeXt $288\times 288$ 92% (71%) 0.61 94% (77%) 8.20
SwinV2 $384\times 384$ 93% (72%) 0.53 95% (80%) 12.69
SwinV2(Perceiver) 0.36 76% (53%) 22.65
pooled ResNet50 $224\times 224$ 89% (68%) 0.12 48% (23%) 31.64
@@
ConvNeXt $288\times 288$ 92% (71%) 0.19 44% (20%) 31.67
SwinV2 $384\times 384$ 93% (72%) 0.16 47% (22%) 49.04
A.2 Control Strength vs. Guidance Scale
@@
In Fig. 9, we present several original images alongside their respective reconstructions, generated using the model trained with the ConvNeXt backbone. These samples were produced with varying levels of control strength and guidance scale. Since both significantly influences the output, we evaluated the model’s performance across different scale settings, again using the ConvNeXt-based model. To quantify the quality of the generated images, we calculated the Top-1 classification match as well as the cosine similarity, illustrated in Fig. 10 and 11. Both correlate strongly and show the same goog performing ranges. This suggests that the unpooled feature map requires minimal text guidance, whereas for the pooled variant, guidance is of substantial importance, as natural looking images cannot be generated without its influence.
(a) Reconstructions for class Magpie
(b) Reconstructions for class Kuvasz
@@
Figure 9: Unpooled and pooled reconstructions with different control strength and guidance scales from ConvNeXt feature maps for class magpie and kuvasz. The cosine similarity between the original feature map and that of the reconstruction is noted at the bottom edge of the images. A low control strength already results in a sufficient reconstruction. While increasing the guidance scale in unpooled samples results in higher color saturation and more distorted object shapes, the same increase improves object fidelity in pooled feature maps. This trend is also roughly reflected in the corresponding cosine similarity.
(a) ResNet50
(b) ConvNeXt
(c) SwinV2
Figure 10: Impact of control strength and guidance scale on generation quality (top-1 accuracy). Top-1 classification accuracy for ResNet50, ConvNeXt and SwinV2 backbone with and without pooling. One image per class was reconstructed for each combination. Higher values indicate better performance. The unpooled model does not seem to rely on guidance as much as the pooled model.
(a) ResNet50
(b) ConvNeXt
(c) SwinV2
Figure 11: Impact of control strength and guidance scale on generation quality (cosine-sim). Cosine similarity (x100) for ResNet50, ConvNeXt and SwinV2 backbone with and without pooling. One image per class was reconstructed for each combination. Higher values indicate better performance. The unpooled model does not seem to rely on guidance as much as the pooled model.
A.3 Variability across reconstructions
As the model we use for reconstruction is a probabilistic model, we show the variance of the generations qualitatively when different seeds are used. This can be seen in Figure 12. The observed variability supports the necessity of a probabilistic rather than a deterministic mapping for the ill-posed problem of mapping from feature space to input space.
Figure 12: Reconstruction variations of feature maps from ResNet50, ConvNeXt, and SwinV2, each unpooled and pooled for the classes volcano, strawberry, and flamingo. It is noticeable that the unpooled reconstructions mostly resemble each other; only ResNet50 displays noticeable differences. However, the reconstructions on pooled feature maps are different. Here, the images appear very diverse, even though they are semantically similar. For example, the volcano class in ResNet50 pooled shows a type of underwater landscape consistently, while the pooled images in ConvNext all show red bushes. Consistency between the reconstructions of the pooled feature maps can also be seen in the other classes, although these are never as strong as those of the unpooled feature maps.
A.4 Generation on unpooled vs. pooled feature maps
@@ A.3 Variability across reconstructions
A.5 Comparison to RCDM
Approach In this section, we provide a direct comparison to the RCDM approach Bordes et al. (2022) as our closest competitor. The RCDM approach conditions on pooled feature representations and can therefore be most meaningfully compared to FeatInv conditioned on pooled feature maps, which receives the same input. We stress at this point, that this experiment facilitates a comparison of different approaches to train diffusion models conditioned on pooled feature maps but does not touch the central novelty of the submission, i.e., conditioning on unpooled feature maps. Whereas RCDM trains a conditional diffusion model from scratch, FeatInv ’s ControlNet approach leverages a pretrained diffusion model.
ResNet50 We present the results for a RCDM model Bordes et al. (2022) conditioned on pooled ResNet50 feature maps, which produces samples at a resolution of 128×128 pixels. Here, we leverage model weights provided by the original authors. To evaluate the quality of the generated images, we computed the same metrics as for the FeatInv models presented in the main text, as shown in Table 4. The results are slightly non-trivial to compare between RCDM and FeatInv as the output resolutions differ (128×128 vs. 256×256), which leaves the option of either evaluating RCDM outputs at their native resolution or scaling it up to 224×224. Irrespective of the precise evaluation procedure, RCDM shows slightly higher cosine similarities (0.16-0.29 vs. 0.12) and higher FID scores (13.0-28.3 vs. 31.6) than pooled FeatInv models. However, these values remain lower than those achieved by our approach using unpooled feature maps.
ConvNeXt We also trained an RCDM model ourselves based on ConvNeXt as ResNet50 showed the worst results across all investigated models in the main text. We therefore used the official RCDM code repository to train a RCDM model conditioned on pooled ConvNeXt feature maps, as before on a resolution 128×128. After 95 hours of training on four L40 GPUs with a batch size of eight, the model had reached 520,000 steps. It performs worse than our FeatInv  equivalents which shows that it either did not fully convergence, indicating that further training would be needed or is not capable of recreating matching samples based on the ConvNeXt features. Our FeatInv  models, which were trained using a smaller batch size of two for less time, saw less images after 1,920,000 steps and performed better. This demonstrates the benefit of using a pre-trained diffusion model rather than training one from scratch.
Conclusion While RCDM is technically capable of scaling to higher resolutions than 128x128, doing so requires long training times. In our experiments, we trained RCDM at half the resolution of our FeatInv  models and for a longer duration, yet its performance still lagged behind FeatInv . In contrast, FeatInv  could be adapted to even higher input resolutions as it leverages a pretrained diffusion model. Results are difficult to comparable directly due to different input resolutions. RCDM seems slightly superior in the case of ResNet50, but clearly inferior in the case of ConvNeXt. In any case, the RCDM approach is not able to deliver models conditioned on unpooled feature maps, which yield clearly superior reconstruction quality and better FID scores. This clearly positions FeatInv as the more versatile approach.
Table 4: Comparison between FeatInv and RCDM: We computed cosine similarity between feature maps of generated and original images. Since there is a significant gap between the dimensions of the original and generated samples, we also examined the results when both were scaled to the same low input size (resulting in 4x4 feature maps), which increased the similarity. Classification accuracy and FID score likewise improved when both inputs were resized to the lower resolution. The ConvNeXt RCDM does not reach the results of our FeatInv equivalent after seeing more than the same amount of images in training.
Model input cosine-sim top5(1) match FID
RCDM ResNet50 $128\times 128$ 0.29 69% (42%) 13.02
ResNet50 $224\times 224$ 0.16 65% (37%) 28.26
ConvNeXt $128\times 128$ 0.12 26% (10%) 37.20
ConvNeXt $224\times 224$ 0.14 26% (10%) 57.46
FeatInv ResNet50 $224\times 224$ 0.46 91% (70%) 11.49
ConvNeXt $288\times 288$ 0.61 94% (77%) 8.20
ResNet50 pooled $224\times 224$ 0.12 48% (23%) 31.64
ConvNeXt pooled $288\times 288$ 0.19 44% (20%) 31.67
Appendix B Conditioning on intermediate feature layers
To condition on intermediate feature layers, we adjusted the input and hidden dimensions of our input encoder to the different dimensions in the model. Unlike the model on the last feature map, each of these models was trained for only one epoch. We observed that training on the last feature maps from early ConvNeXt stages in particular led to good reconstruction results much faster than training on features from the last layer. For these models, we also reconstructed 10 validation images per class (see Figure 14) and calculated cosine similarity, FID and matching accuracies (see Figure 4). For the reconstruction we provide the last feature map from the stage the model was trained on.
In Figure 14, it can be seen that the models that reconstruct earlier feature maps work more accurately and display details better (e.g., watermark text). Colors and edges within the image are also displayed better.
Figure 14: Comparison of reconstructions of models trained on the last feature maps from different ConvNeXt stages with dimensions (128, 72, 72), (256, 36, 36), and (512, 18, 18). The cosine similarity of the original feature map and that of the reconstruction is noted at the bottom edge of the images.
Appendix C Reconstruction of corrupted images
We used FeatInv to reconstruct corrupted images as described in Section 4.1. Both the FID score (see Figure 15) and cosine similarity (see Figure 16) suggest that FeatInv can only represent these manipulation up to a certain level. We show the top5(1) accuracy matches for the corrupted images and their reconstructions (see Figure 17(18)). In addition we present some examples of what the reconstruction of the different categories with FeatInv looks like (see Figures 19, 20, 21, 22 and 23).
Figure 15: Comparison of the FID scores of the ImageNet-C categories. While a slight corruption of the images, as in level 1, only leads to a slight increase in the FID score, a high corruption has a significant impact on the score. The higher the corruption, the more difficult it is for FeatInv to reconstruct the corrupted images. It should be noted that the model was not trained on any of these corruptions. Noise and contrast in particular have a high negative impact on the scores and cannot be reproduced well by FeatInv.
Figure 16: Comparison of the cosine similarity of the feature maps of the corrupted input images and their reconstructed images. Both minor (level 1) and major corruption (level 5) have a strong influence on the cosine similarity of the final feature maps and cause them to drop. It should be noted again that the model was not trained on any of these corruptions. Unlike the FID scores of the images in these different categories, there does not seem to be any category in which the change differs significantly from that of the other categories.
Figure 17: Comparison of the top5 accuracy matches of the ImageNet-C categories. While a slight corruption such as in level 1 achieves almost the same value in top5 accuracy matching as the unaltered images, a stronger corruption such as in level 5 has a much greater negative impact on the matching. Compared to the cosine similarity in Figure 16, the matching works unexpectedly well, which is most likely due to the fact that the model assigns the classes of the two images, one of which was reconstructed by featinv and may not contain the complete corruption, to the same class. The result is different when there is a large difference in the reconstruction.
Figure 18: Comparison of the top1 accuracy matches of the ImageNet-C categories. As expected, the results of top1 accuracy matching correspond to those of top5 accuracy matching. As with the unaltered images, the values for top1 matching are lower than those for top5 matching.
Figure 19: Reconstructed corrupted images: Noise (gaussian noise, shot noise, and impulse noise). In each category, the original corrupted images can be seen at the top and the reconstructed images below. On the left is level 1, i.e., weak corruption, and on the right is level 5, i.e., strong corruption. It can be seen that FeatInv is not very good at reconstructing detailed noise, and the images lose detail. In particular, the colors of the noise are lost.
Figure 20: Reconstructed corrupted images: Blur (defocus blur, glass blur, motion blur, and zoom blur). 0 In each category, the original corrupted images are shown above and the reconstructed images below. On the left is level 1, i.e., weak corruption, and on the right is level 5, i.e., strong corruption. It can be seen that FeatInv is quite capable of reconstructing the images. Both levels 1 and 5 show good results.
Figure 21: Reconstructed corrupted images: Weather (frost, snow, fog, and brightness.) In each category, the original corrupted images are shown above and the reconstructed images below. On the left is level 1, i.e., weak corruption, and on the right is level 5, i.e., severe corruption. It can be seen that FeatInv is quite capable of reconstructing the images. However, there are images where the reconstruction does not work as well, such as the second image, the flying bird. In all categories except fog, the reconstruction differs greatly from the original corrupted image.
Figure 22: Reconstructed corrupted images: Digital (contrast, elastic transform, pixelate, and jpeg compression.) In each category, the original corrupted images are shown above and the reconstructed images below. On the left is level 1, i.e., weak corruption, and on the right is level 5, i.e., severe corruption. It can be seen that FeatInv is quite capable of reconstructing the images in the last three categories well. However, this does not work for contrast. Especially at level 5, the reconstructed images differ significantly from the original corrupted image. The main object is more emphasized than in the image with low contrast.
Figure 23: Reconstructed corrupted images: Extra (speckle noise, spatter, gaussian blur and saturate.) In each category, the original corrupted images are shown above and the reconstructed images below. On the left is level 1, i.e. weak corruption, and on the right is level 5, i.e. strong corruption. Except for the first category, speckle noise, FeatInv is able to reconstruct the images well. In the case of spatter, the dots are reconstructed, but some are located differently or are fewer in number. In the case of speckle noise, similar to the reconstructions in Figure 19, the noise is not well represented in the reconstructed images.
Appendix D Details on concept steering